OpenStudy (anonymous):

find the vertex the line of symmetry the maximum or minimum value of the quadratic function and graph function. f(x)=-2x^2+2x+7 I have the x-coordinate vertex is (-2,7)

OpenStudy (anonymous):

f(x) = -2x^2 + 2x + 7 the vertex is at x = -b/2a = -2/-4 = 1/2 Huh?

OpenStudy (anonymous):

(-2, 7) is not a solution.

OpenStudy (anonymous):

that answer was given already.

OpenStudy (anonymous):

You can try it for yourself -2(4) + 2(-2) + 7 = -5